OpenStudy (anonymous):

solve each system by substitution Please show how you got it and also how you found Y... Thanks :) 4x+2y=10 x-y=13

OpenStudy (anonymous):

x = 13 + y 4(13 + y) + 2y = 10 y = -7 x = 13 - 7 x = 6 (6, -7)

OpenStudy (anonymous):

Thankks you